OpenStudy (anonymous):

Solve each equation 19. 3(2x-5)=x+3x+7 20. 2y-6y=y+31-49

OpenStudy (anonymous):

simplify first, then move like terms to the same side, and solve for the variable

OpenStudy (anonymous):

can you show me how to do it with numbers please so i can have sumn visual?

OpenStudy (anonymous):

so for 19: 6x-15=4x+7 2x=22 x=11

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yess thanks

OpenStudy (anonymous):

welcome (:

OpenStudy (anonymous):

can you help with 20?

OpenStudy (anonymous):

2y-6y=y+31-49 -4y=y-18 -3y=18 y=-6

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np